- Calculate I (x-1)² -oo (x² + 4x+8) ² dx
Answer #1

f (x-17² dx ao C2392+8) tix Let 6x) = 6L-132 of 9(x) = (x?+4x+8)2 P & as As gle) can be differentiate in two repeated roots kdx =3 da J8x²+42+8) 8 Ja²+uat8 (ar 13 l dx llo: (x 80x²+2.22 +2²+4 sl (c) & Bl de ily 8 + 2)² + 4 x (N) oo cool at2 = a Let sPutting this value ean. lid we get 5x+34 + 13 tan(at2) +Clu 8(x² + 4x+8) 16 I ator lots oder nut as a strollib sooroo OOR Put